OpenStudy (anonymous):

Daria can wash and detail 3 cars in 2 hours. Larry can wash and detail the same 3 cars in 1.5 hours. About how long will it take to wash and detail the 3 cars Daria and Larry worked together?

OpenStudy (anonymous):

So Daria washes 1.5 cars per hour, and Larry washes 2 cars per hour. Assuming they work together perfectly, they can wash 3.5 cars per hour. Put differently, together they wash 1/3.5 = 2/7 cars per hour. So to wash 3 cars, it will take 3 * 2/7 = 6/7 of an hour. This works out to be about 51 minutes. 6/7 hr * 60 min/hr = 51 min.

OpenStudy (anonymous):

wow thank you so much